OpenStudy (anonymous):

Combine: square root 90 - square root 10

OpenStudy (anonymous):

\[\sqrt{90}-\sqrt{10}\]

OpenStudy (anonymous):

\[=\sqrt{80}\]

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

not true that \(\sqrt{25}-\sqrt{9}=\sqrt{16}\) is it? that would be like saying \(5-3=4\)

OpenStudy (anonymous):

first rewrite \(\sqrt{90}=\sqrt{9\times 10}=\sqrt{9}\sqrt{10}=3\sqrt{10}\) then you have \[3\sqrt{10}-\sqrt{10}\] and combine like terms

OpenStudy (anonymous):

I would combine the 10's

OpenStudy (anonymous):



OpenStudy (anonymous):

did I do it right

OpenStudy (anonymous):

not quite you cannot add or subtract radicals that way \[3x-x=2x\] and so \[3\sqrt{10}-\sqrt{10}=2\sqrt{10}\]